Transformations at Work Solve the IVPs in Problem using Laplace transforms. y + y = cos 3t; y(0) = 1, y(0) = -||

Please substitute cos 3t with cos 2t instead. Thank you
Answer #1

Given IvP is y+gy - Cos 27 410)=1 440=-1 Tere laplace transform Lag(t)} +9 Lsycts} = 24 coszt} 52715) - SY10) –y:o) + 9 YCSS Iss rss s2+9 (S2+4) (5²9) 5274 S .., = 5} r44) 65478] 54352#x}-324 5²+g Š cosat - s dosat also s 1 = Cosat Ik ;] 1{shtg}